solve: -2/3x -5/6=3/8
Answer by CharlesG2(834)

solve: -2/3x -5/6=3/8

(-2/3)x - 5/6 = 3/8
(-2/3)x = 9/24 + 20/24 (3/8 = 9/24, 5/6 = 20/24)
(-2/3)x = 29/24
x = -3/2 * 29/24 (reciprocal of -2/3 is -3/2)
x = -87/48 = -29/16
check:
-2/3 * -29/16 - 5/6 = 3/8
58/48 - 40/48 = 18/48 (5/6 = 40/48, 3/8 = 18/48)
18/48 = 18/48